# QUASAR: A Universal Autonomous System for Atomistic Simulation and a Benchmark of Its Capabilities

### v1

@@ 5 Summary and Outlook
Through our Tier III benchmarking, QUASAR demonstrates that LLMs, when properly orchestrated, are capable of executing sophisticated computational chemistry tasks with a level of scientific rigor comparable to that of human researchers. This work serves as a proof of concept that modern AI has the potential to automate the vast majority of standard computational workloads, freeing the scientific process from the bottlenecks of manual execution. This positions QUASAR as a prototype for human–AI collaborative workflows in computational chemistry, where AI handles routine orchestration while humans provide scientific direction, judgment on ambiguous cases, and validation of results.

### v2

@@ 5 Summary and Outlook
Through our Tier III benchmarking, QUASAR demonstrates that LLMs, when properly orchestrated, are capable of executing sophisticated computational chemistry tasks with a level of scientific rigor comparable to that of human researchers. This work serves as a proof of concept that modern AI has the potential to automate the majority of standard computational workloads, freeing the scientific process from the bottlenecks of manual execution. This positions QUASAR as a prototype for human–AI collaborative workflows in computational chemistry, where AI handles routine orchestration while humans provide scientific direction, judgment on ambiguous cases, and validation of results.